Jamie is doing a survey at her school about whether the students feel the cafeteria food is satisfactory. She records the responses of 150 students and finds that 42% find the food satisfactory. She wonders whether she can claim this is significantly less than half.

A)Describe the sampling distribution. ( Please explain in detail)

B) Find and interpret the 96% confidence interval for the proportion of students who find the cafeteria food satisfactory. (Please explain in detail)

C)What type of hypothesis test should be performed here?

Options: Z-Test | T-Test | None of these. | 1-Prop Z Test

D)Detail your testing process. Include all relevant information such as hypothesis, critical values, conclusions, and interpretations. Assume your significance level is 0.1. ( Please explain in detail)
Answer #1

Ans:

hypothised proportion=0.5

a)sampling distribution of sample proportion:

mean=p=0.5

standard deviation=sqrt(0.5*(1-0.5)/150)=0.0408

b)critical z value(2 tailed)=2.054

96% confidence interval for the proportion

=0.42+/-2.054*sqrt(0.42*(1-0.42)/150)

=0.42+/-0.083

=(0.337, 0.503)

c) 1-Prop Z Test

d)

Test statistic:

z=(0.42-0.5)/sqrt(0.5*(1-0.5)/150)

z=-1.96

critical z value=normsinv(0.1)=-1.282

As test statistic is less than -1.282,we reject the null hypothesis and conclude that proportion of students who feel the cafeteria food is satisfactory is significantly less than half.